A circular pool measures 10 feet across. One cubic yard of concrete is to be used to create a circular border of uniform width around the pool. If the border is to have a depth of 3 inches, how wide will the border be? Use 3.14 to approximate pi. Express your solution rounded to two decimal places. (1 cubic yard = 27 cubic feet)
I don't know how to do this problem. I simply followed a sample found on this website.
Let d = depth of the border
Let w = width of border to be found
Let V = volume of pool
Let r = radius of pool
V = d•pi(w + r)^2 - pi (25•d) = 27
V = 3•pi(w + 5)^2 - pi(25•3) = 27
V = 3•pi(w + 5)^2 - 75•pi = 27
Subtract 27 on both sides.
Set V = 0 and solve for w.
0 = 3•pi(w + 5)^2 - 75•pi - 27
After solving for w, I get two approximations:
-10.27871 and 0.278711
Of course, I must reject the negative value for w because distance must be positive.
The book's answer for the width of the border is 2.71 feet wide.
What am I doing wrong?
Thank you.




Found 2 solutions by ikleyn, htmentor:
Answer by ikleyn(52781)

.

Your error is that you use wrong value for the thickness  " d "  of the circular walkway.

According to the problem,  it is  3 inches (sic !),  which is  0.25  of the foot.


*************************************

        1 foot = 12 inches

*************************************


So the correct value for  d  is   d = 0.25 ft;   NOT   d = 3 ft.



        They caught you in the trap,  using different measurement units . . .



Happy calculations and re-calculations  ( ! )



Answer by htmentor(1343)

It looks like the only place you went wrong was you did not convert the depth in inches to feet. You should have d = 3/12 = 1/4 ft. You are mixing units in your equations. Repeat with d = 1/4 and you should get the right answer.